OpenStudy (anonymous):

Solve for x. |2x + 6| - 4 = 20 x = 4 and x = 0 x = -15 and x = 9 x = -4 and x = -2 x = -3 and x = 1 no solutions

OpenStudy (anonymous):

method or answer?

OpenStudy (anonymous):

only reason i ask is because you have posted a bunch of these and the method is always the same

OpenStudy (anonymous):

answer

OpenStudy (ash2326):

@moody22 Please try, I can see that you are posting similar questions and just looking for answer. Please try to understand the fundamentals, it'll help you in the long run. Thanks

OpenStudy (anonymous):

then no need to post here where you might get a mistaken answer just do this http://www.wolframalpha.com/input/?i= |2x+%2B+6|+-+4+%3D+20+

OpenStudy (anonymous):

like the one above

OpenStudy (anonymous):

just copy and paste the link i sent you will get all the answers you want, and no mistakes if, on the other hand , you want the method, then ask here

OpenStudy (anonymous):

really? lets check \(x=0\) \(|2x + 6| - 4 = 20\) \[|2\times 0+6|-4=20\] \[|0+6|-4=20\] \[6-4=20\] \[2=20\]